# Quantum Dynamic Programming

### v1

@@ Abstract
We introduce a quantum extension of dynamic programming, a fundamental computational method for efficiently solving recursive problems using memory. Our innovation lies in showing how to coherently generate unitaries of recursion steps using memorized intermediate quantum states. We find that quantum dynamic programming yields an exponential reduction in circuit depth for a large class of fixed-point quantum recursions, including a known recursive variant of the Grover’s search. Additionally, we apply quantum dynamic programming to a recently proposed double-bracket quantum algorithm for diagonalization to obtain a new protocol for obliviously preparing a quantum state in its Schmidt basis, providing a potential pathway for revealing entanglement structures of unknown quantum states.
Introduction.—Access to placeholder memory can drastically reduce computation runtime. For example, there is an exponential difference in computing the classic Fibonacci sequence with or without placeholder memory. Should one attempt it directly from the recursive definition
$F(n)=F(n-1)+F(n-2),\qquad n\geq 2,$ (1)
starting with $F(0)=0$ and $F(1)=1$, it is quickly realized that in order to compute $F(N)$ naively, we need to compute $F(N-2)$ twice because it appears in the expression for $F(N)$ and also for $F(N-1)$. In turn computing $F(N-2)$ requires two computations of $F(N-4)$, and so on. This leads to $\mathcal{O}(2^{N})$ additions to arrive at $F(N)$. Fortunately, one can mitigate this computational overhead by the elegant solution of memoization [1], i.e. using classical memory to store values of intermediate states $F(n)$, and retrieving $F(n)$ whenever necessary. The program then contains an explicit section (e.g. a memory) that changes with time and $F(N)$ is computed in $\mathcal{O}(N)$ time. Such strategies are known as dynamic programming [2, 3], and they utilize memory calls of the appropriate quantity to yield vastly shorter runtime than re-computations.
Memoization is more difficult in quantum computing than for classical computers because the readout of a quantum state $\rho$ renders it useless for future ‘retrieval’. As opposed to classical computing, this cannot be circumvented by the cloning of information [4]. Nonetheless, we will show that quantum dynamic programming, i.e. involving memoization in quantum algorithms, remains strategically the right choice and can lead to drastic circuit depth reductions, en par with the classic example of computing Fibonacci numbers.
Due to the no-cloning theorem, quantum memory is ‘consumed’ upon each ‘retrieval’, necessitating the preparation of multiple independent copies of an intermediate state. Hence, to achieve reductions in circuit depth, one needs circuits with a larger width. The crucial difference to the precomputation approaches [5, 6] emerges here; our memory states are not given, but we prepare them as a result of previous recursion steps. The localized structure of the arising circuits alleviates the difficulties associated with the large width requirement. Quantum memoization operates by partitioning a large computational task into smaller and independent state preparation circuits, much simpler than the full task and capable of being executed in parallel. Furthermore, the quantum recursions of interest exhibit strongly attractive fixed-points, i.e. the task at hand includes an intrinsic stabilizing effect so that, if sufficiently small, errors of the intermediate states do not wreck the quantum computation. The local modularity and intrinsic stabilization make such recursions a natural use-case for distributed quantum computing [7, 8, 9], when implemented using quantum dynamic programming.
Any protocol must conclude within the coherence time scales of the quantum processor; therefore, among circuits achieving the same task, it is strategic to select those with minimized depth rather than width. Quantum dynamic programming radically alters the circuit structure to achieve just that, i.e. the key characteristic — circuit depth — gets shortened through quantum memoization. We show that this underexplored design principle can enable successful quantum computations where ordinary proposals would fail.
Quantum recursion steps consist of unitary transformations $\hat{U}^{(\rho_{n})}$ that explicitly depend on past states $\rho_{n}$,
$\rho_{n+1}=\hat{U}^{(\rho_{n})}\rho_{n}\left(\hat{U}^{(\rho_{n})}\right)^{\dagger}\ .$ (2)
Importantly, $\rho_{n}$ may be unknown, so in each recursion step $n$, the transformation must be implemented ‘obliviously’. Thus, Eq. (2) entails a key challenge for using memoization in quantum algorithms: How can an intermediate state $\rho_{n}$ feed back into the ‘knobs’ of the quantum computer and determine the subsequent gate applications? In other words, how can we use quantum data $\rho_{n}$ as a quantum instruction [10] for $\hat{U}^{(\rho)}$?
@@
In this manuscript, we systematically address this challenge for a class of recursion unitaries $\hat{U}^{(\rho)}$ by employing quantum memory-calls. Specifically, we apply it to i) a known quantum algorithm, namely recursive Grover’s search [11], and ii) a dynamic implementation of recently proposed quantum recursions for diagonalization [12], leading to a quantum algorithm for transforming a state into its Schmidt basis without knowledge of the state itself. With this, we establish that a quantum-mechanical version of dynamic programming is not only a strategic conceptual framework but also highly applicable.
Memory-calls in quantum recursions.—Suppose we desire to implement a unitary that depends on a state $\rho_{n}$, as in Eq. (2), but we do not know $\rho_{n}$ and only have access to copies of $\rho_{n}$. To characterize such unitaries, let $\hat{\mathcal{N}}$ be a map from Hermitian quantum instruction states $\rho$ to Hermitian operators $\hat{\mathcal{N}}(\rho)$. Thus, $\hat{\mathcal{N}}$ is Hermitian-preserving and $e^{i\hat{\mathcal{N}}(\rho)}$ is unitary as required. We say that a quantum recursion contains a single memory-call, if it is of the form
$\hat{U}^{(\hat{\mathcal{N}},\rho)}=\hat{V}_{2}e^{i\hat{\mathcal{N}}(\rho)}\hat{V}_{1}\ .$ (3)
Here $\hat{V}_{1},\hat{V}_{2}$ are static unitaries, independent of the instruction state $\rho$. That state can vary dynamically so the memory-call unitary $e^{i\hat{\mathcal{N}}(\rho)}$ is dynamic too.
Quantum programming with memory-calls consists of choosing $\hat{\mathcal{N}}$ appropriately. For example $\hat{\mathcal{N}}(\rho)=t\rho$ corresponds to density matrix exponentiation [13, 14] and enables a framework for universal quantum computation. Another example is $\hat{\mathcal{N}}(\rho)=[\hat{D},\rho]$, the commutator of the input with a fixed operator $\hat{D}$ which facilitates diagonalization [12], as we will see in our example algorithm. Eq. (3) naturally generalizes to a recursion with multiple memory-calls
$\hat{U}^{(\{\hat{\mathcal{N}}\},\rho)}=\hat{V}_{L}e^{i\hat{\mathcal{N}}_{L}(\rho)}\hat{V}_{L-1}e^{i\hat{\mathcal{N}}_{L-1}(\rho)}\ldots\hat{V}_{1}e^{i\hat{\mathcal{N}}_{1}(\rho)}\hat{V}_{0}\ .$ (4)
We will next discuss a recursive variant of Grover’s algorithm [15] whose recursion steps are exactly in this form.
Unfolding quantum recursions—The nested fixed-point Grover search [11] has the recursive structure $|{\psi_{n+1}}\rangle=\hat{\Gamma}^{(\psi_{n})}_{L}|{\psi_{n}}\rangle$, where each step consists of alternating reflections
$\hat{\Gamma}^{(\psi)}_{L}=\displaystyle\prod_{i=1}^{L}\hat{E}^{(\psi)}_{\alpha_{i}}\hat{E}^{(\tau)}_{\beta_{i}}.$ (5)
Here $\hat{E}^{(\sigma)}_{s}=e^{-is\sigma}=\mathmybb{1}-(1-e^{-is})|{\sigma}\rangle\!\langle{\sigma}|$ are reflection unitaries and $\tau$ is the pure target state which the algorithm aims to output. In the recursion, Eq. (5) depends dynamically on any intermediate state $\psi_{n}$. The set of predetermined angles $\{\alpha_{i},\beta_{i}\}_{i=1}^{L}$ optimizes the procedure and if we set $\hat{\mathcal{N}}_{i}(\rho)=-\alpha_{i}\rho$ and $\hat{V}_{i-1}=\hat{E}^{(\tau)}_{\beta_{i}}$, $\hat{V}_{L}=\mathmybb{1}$ then Eq. (5) is in form of Eq. (4).
Ref. [11] proposed to use re-computations to prepare the state $|{\psi_{N}}\rangle=\hat{\Gamma}^{(\psi_{N-1})}_{L}\ldots\hat{\Gamma}^{(\psi_{0})}_{L}|{\psi_{0}}\rangle\approx|{\tau}\rangle$. No placeholder memory beyond the register storing $|{\psi_{0}}\rangle$ is involved; hence, we call it a zero-memory approach. Consider the simplest case $L=1$, when the recursion involves only one memory-call. For the second recursion step we have $|{\psi_{2}}\rangle=\hat{\Gamma}^{(\psi_{1})}_{1}\hat{\Gamma}^{(\psi_{0})}_{1}|{\psi_{0}}\rangle$, with $\hat{\Gamma}^{(\psi_{1})}_{1}$ and $\hat{\Gamma}^{(\psi_{0})}_{1}$ defined in Eq. (5). The reflection around $\psi_{1}$ is not directly available but we assume that we can reflect around $\psi_{0}$ and $\tau$. Crucially $\psi_{1}=\hat{\Gamma}^{(\psi_{0})}_{1}{\psi_{0}}\bigl{(}\hat{\Gamma}^{(\psi_{0})}_{1}\bigr{)}^{\dagger}$ implies
$\hat{E}^{(\psi_{1})}_{\alpha_{1}}=\hat{\Gamma}^{(\psi_{0})}_{1}\hat{E}^{(\psi_{0})}_{\alpha_{1}}\left(\hat{\Gamma}^{(\psi_{0})}_{1}\right)^{\dagger}\ .$ (6)
Consequently, we can implement $\hat{\Gamma}^{(\psi_{1})}_{1}=\hat{E}^{(\psi_{1})}_{\alpha_{1}}\hat{E}^{(\tau)}_{\beta_{1}}$ using the available relfections and obtain $|{\psi_{2}}\rangle$. For further recursion steps we proceed analogously, each time obtaining a sequence of unitaries which involve reflections only around $\psi_{0}$ and $\tau$.
The key step is the covariance Eq. (6). The same approach works for any $L$ when there are multiple memory-calls, and in Sec. II.1 of Supplemental Materials we generalize this zero-memory procedure to multiple memory-call recursion steps where polynomials $f_{k}$ are used in lieu of $\hat{\mathcal{N}}_{k}$ for Eq. (4).
We refer to this strategy of executing quantum recursions as ‘unfolding’ [11, 12]. Its severe drawback is that it needs an exponentially deep circuit – a consequence of applying $\hat{U}^{(\rho_{n-1})}$ and its inverse $L$ times to implement $\hat{U}^{(\rho_{n})}$. Consequently, the circuit depth for evaluating the $n$th recursion step scales as $(2L)^{n}$, in units of the gate complexity needed for an initial recursion step. Unfolding is analogous to a re-computation of intermediate states $\rho_{n}$ and thus is static rather than dynamic.
Memory-usage queries.—We propose to prepare memory-call unitaries $e^{i\hat{\mathcal{N}}(\rho)}$ by dynamically invoking memory. The execution via memoization means that we apply a unitary operation $\mathcal{Q}$ on two registers, the working register containing $\sigma$ and the memory register containing the instruction state $\rho$. After the action of $\mathcal{Q}$, we discard the memory register, while the working state becomes transformed by a dynamic operation instructed by $\rho$. We dub this protocol a memory-usage query, and remark that $\mathcal{Q}$ is oblivious to the states $\rho$ and $\sigma$.
More specifically, we need a way to apply the unitary channel $\hat{\mathbf{E}}^{(\hat{\mathcal{N}},\rho)}$ of $e^{i\hat{\mathcal{N}}(\rho)}$ to a working state $\sigma$:
$\hat{\mathbf{E}}^{(\hat{\mathcal{N}},\rho)}(\sigma)=e^{i\hat{\mathcal{N}}(\rho)}(\sigma)e^{-i\hat{\mathcal{N}}(\rho)}\ .$ (7)
@@
^s$\hat{\mathcal{E}}^{(\hat{\mathcal{N}},\rho)}_{s}(\sigma)=\operatorname{Tr}_{1}\left[e^{-i\hat{N}s}\left(\rho\otimes\sigma\right)e^{i\hat{N}s}\right]\ ,$ (8)
where $\mathcal{Q}=e^{-i\hat{N}s}$ is oblivious, i.e. independent of both the memory and working registers. We ‘consume’ the state $\rho$ as we trace out the memory register. Eq. (8) is a single query because we use one copy of $\rho$; if this procedure is repeated $M$ times with $s=\frac{1}{M}$, i.e. if $M$ queries consuming $M$ copies of $\rho$ are performed, we obtain
^$\hat{\mathcal{E}}^{(\hat{\mathcal{N}},\rho,M)}_{\text{QDP}}\coloneq\left(\hat{\mathcal{E}}^{(\hat{\mathcal{N}},\rho)}_{1/M}\right)^{M}=\hat{\mathbf{E}}^{(\hat{\mathcal{N}},\rho)}+\mathcal{O}\left(1/{M}\right)\ .$ (9)
@@
Quantum dynamic programming.—We define quantum dynamic programming (QDP) to consist of making memory-usage queries on copies of $\rho_{n}$ to approximate a recursion step towards $\rho_{n+1}$. More specifically, consider initializing a quantum program with $\sigma_{0}=\rho_{0}$, and defining the QDP iteration
$\sigma_{n+1}=\hat{\mathcal{E}}^{(\hat{\mathcal{N}},\rho,M)}_{\text{QDP}}(\sigma_{n})\ .$ (10)
@@
From Eq. (9), we see that Eq. (10) approximates a single memory-call recursion (Eq. (3) with $\hat{V}_{1}=\hat{V}_{2}=\mathmybb{1}$) by $M$ memory-usage queries. In general, for recursions with multiple memory-calls, we define the QDP iteration as replacing all memory-calls $e^{i\hat{\mathcal{N}}(\rho)}$ by total $M$ memory-usage queries.
Trading circuit depth for width with QDP.—To prepare one copy of $\sigma_{1}$, we make $M$ memory-usage queries using $\sigma_{0}$ as an instruction. In other words, the first step consumes $M$ root state copies. Likewise, to prepare one copy of $\sigma_{2}$, we need $M^{2}$ copies of $\sigma_{0}$; thus, the iteration of Eq. (10) consumes $M^{n}$ copies of the root state. Although the circuit width grows exponentially with $n$, the circuit depth remains polynomial: the preparation of multiple copies of instruction states is done in parallel, and the maximum depth is equal to performing $nM$ memory-usage queries. In comparison, unfolding requires a circuit depth exponential in $n$ and has a constant width.
Before declaring an exponential depth reduction over unfolding, however, we must ensure that each memory-call approximation in QDP is sufficiently accurate to prepare the final state $\sigma_{N}$ within the same error threshold as unfolding. We next discuss the number of memory-usage queries $M$ needed for that.
Exponential depth reduction with QDP.—Aiming for an exact solution of quantum recursions is unrealistic because of inaccuracies prevalent in quantum circuit compilation, e.g. due to a finite $M$ in Eq. (9), among others. We next discuss the accuracy of QDP when allowing for errors in the preparation of the root state $\sigma_{0}$ and deviations from the exact recursion steps.
By Eq. (9), $M=\mathcal{O}(1/\epsilon)$ suffices for $\|\sigma_{1}-\rho_{1}\|_{1}\leq\mathcal{O}(\epsilon)$. The next step might have an amplified deviation because the second recursion step $\sigma_{2}=\hat{\mathcal{E}}^{(\hat{\mathcal{N}},\sigma_{1},M)}_{\text{QDP}}(\sigma_{1})$ is instructed by $\sigma_{1}$ and not $\rho_{1}$. Indeed, triangle inequality gives $\|\sigma_{2}-\rho_{2}\|_{1}\leq\mathcal{O}(M\epsilon)$; see Sec. II of Supplemental Materials for the full analysis. We shall next characterize a sufficient condition for errors on quantum instructions not to propagate destructively, indicating that this upper bound need not be realized.
It turns out that if the first few steps are sufficiently accurate, implementing the next steps benefits from the stabilization property of the recursions itself. Indeed, for many fixed-point recursions (e.g. by use of the Polyak-Łojasiewicz inequality for gradient descent iterations [17] or by asymptotic stability analysis of time-discrete dynamical systems [18]) exponential convergence can be obtained.
For a quantum recursion, exponential convergence would read
$\|\rho_{N}-\rho_{\infty}\|_{1}\leq\alpha^{N}\|\rho_{0}-\rho_{\infty}\|_{1}\ ,$ (11)
for some $\alpha<1$. However, we must assume that i) a stable fixed-point $\rho_{\infty}$ is reached for $N\rightarrow\infty$, ii) it is ‘spectrally’ unique, i.e. any two root states of the recursion $\rho_{0}$ and $\rho_{0}^{\prime}$ with the same spectrum have the same fixed-point $\rho_{\infty}=\rho_{\infty}^{\prime}$, and finally iii) the fixed-point attracts all states sufficiently strongly to tolerate small implementation errors. The latter property is made explicit in Sec. III of Supplemental Materials and plays the role of deciding the convergence rate similar to $\alpha$ in Eq. (11).
We say that a quantum recursion satisfying conditions i,ii,iii) has fast spectral convergence. Conversely, the absence of iii) renders the recursion unstable, suggesting its lack of physical relevance; protocols achievable in nature must be resilient to small perturbations. In such cases quantum computation cannot be expected to be both efficient and successful, potentially necessitating infinite precision and thus infinite resources for convergence. Implementing recursions with unfolding is subject to similar stability requirements. Thus, we should focus on recursions with fast spectral convergence; otherwise, the computational task is ill-conditioned.
If a recursion exhibits fast spectral convergence, QDP delivers an exponential depth reduction compared to unfolding:
Theorem 1 (QDP, non-technical).
Suppose that the quantum recursion satisfies fast spectral convergence. If we implement $N$ recursion steps of the form Eq. (4), by replacing each memory-call by $\mathcal{O}(N\epsilon^{-1})$ memory-usage queries, the final state $\sigma_{N}$, such that
$\|\sigma_{N}-\rho_{N}\|_{1}\leq\epsilon,$ (12)

### v2

@@ Abstract
We introduce a quantum extension of dynamic programming, a fundamental computational method that efficiently solves recursive problems using memory. Our innovation lies in showing how to coherently generate recursion step unitaries by using memorized intermediate quantum states. Quantum dynamic programming achieves an exponential reduction in circuit depth for a broad class of fixed-point quantum recursions, though this comes at the cost of increased circuit width. Interestingly, the trade-off becomes more favourable when the initial state is pure. By hybridizing our approach with a conventional memoryless one, we can flexibly balance circuit depth and width to optimize performance on quantum devices with fixed hardware constraints. Finally, we showcase applications of quantum dynamic programming to several quantum recursions, including a variant of Grover’s search, quantum imaginary-time evolution, and a new protocol for obliviously preparing a quantum state in its Schmidt basis.
Introduction.—Storing intermediate results in memory can drastically reduce computation runtime. Computing the Fibonacci sequence provides a perfect example. Direct calculation of the $N$th element using the recursive definition $F(n)=F(n-1)+F(n-2)$, starting with $F(0)=0$ and $F(1)=1$ and without storing intermediate results, requires calculating $F(N-2)$ twice—once for $F(N)$ and again for $F(N-1)$. In turn, computing $F(N-2)$ requires two computations of $F(N-4)$, and so on. This repetition causes the computation time to grow exponentially with $N$. On the other hand, if we store intermediate values of $F(n)$ in memory and retrieve them whenever necessary, $F(N)$ can be computed in time linear to $N$, achieving an exponential speedup. This technique, called memoization [1], is a form of dynamic programming [2, 3] that utilizes a small amount of memory to yield vastly shorter runtime by avoiding re-computations.
We introduce quantum dynamic programming (QDP) as a technique to utilize memoization in quantum computation. The quantum analogue of computing the Fibonacci sequence would be solving quantum recursions, which we define as
$|{\psi_{n}}\rangle\mapsto|{\psi_{n+1}}\rangle=\hat{U}^{[\psi_{n}]}|{\psi_{n}}\rangle\ ,$ (1)
a sequence of quantum states $\{|{\psi_{n}}\rangle\}_{n}$ related by unitary operators $\hat{U}^{[\psi_{n}]}$ that explicitly depend on the past states $|{\psi_{n}}\rangle$. Such recursions appear in important quantum algorithms, including nested fixed-point search [4, 5] and double-bracket quantum algorithms [6, 7, 8, 9].
There are various approaches to solving Eq. (1), yet a uniquely quantum challenge underlies them all. Unlike classical computing, reading out a quantum state $|{\psi}\rangle$ requires many copies, and this cannot be circumvented [10] as dictated by the no-cloning theorem [11]. If one compiles the recursion step $\hat{U}^{[\psi_{n}]}$ by learning each intermediate state $|{\psi_{n}}\rangle$ through, e.g. tomography [12], each step must be repeated many times before learning how to compile $\hat{U}^{[\psi_{n}]}$. An alternative approach, termed unfolding, is analogous to classical computation without memoization. It makes multiple queries to an operation preparing $|{\psi_{n}}\rangle$ in order to compile $\hat{U}^{[\psi_{n}]}$. However, since $|{\psi_{n}}\rangle$ is not learnt, operations preparing it must be compiled using queries to the initial step $\hat{U}^{[\psi_{0}]}$, which is assumed to be given. This leads to the circuit depth exponential in $N$ for preparing $|{\psi_{N}}\rangle$. We expound on this scaling later in this Letter with an example.
Our approach to QDP belongs to a lineage of studies on circuits instructed by quantum states [13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23], where desired operations are implemented by injecting quantum instruction states that encode the operation, rather than by compiling circuits based on classical information. This paradigm offers several advantages. First, the circuit can be implemented obliviously to the quantum instructions, i.e. without learning them, while requiring significantly fewer samples compared to tomography-based learning-and-compiling methods [16, 22]. Furthermore, the circuits instructed by quantum states are universal [16], enabling all queries to $\hat{U}^{[\psi_{0}]}$ (required in the unfolding approach) to be replaced by the injection of $|{\psi_{0}}\rangle$.
The main breakthrough of our work is the following observation: instead of unfolding the recursion (i.e. compiling $\hat{U}^{[\psi_{n}]}$ with exponentially many queries to $\hat{U}^{[\psi_{0}]}$), we can dynamically evolve memory states from $|{\psi_{0}}\rangle$ to $|{\psi_{n}}\rangle$ in parallel and directly implement $\hat{U}^{[\psi_{n}]}$ by injecting these memory states. In essence, QDP ‘folds up’ the unfolded recursion step $\hat{U}^{[\psi_{n}]}$ into the memory state as it evolves to $|{\psi_{n}}\rangle$. By doing so, we achieve an exponential circuit depth reduction compared to the unfolding approach, as formalized in Theorem 1, en par with the classic example of computing Fibonacci numbers.
However, the same no-cloning challenge prevails QDP. Many copies of each intermediate state $|{\psi_{n}}\rangle$ are required, necessitating circuits with greater width. The crucial distinction from precomputation approaches [14, 24] emerges here: our memory states are not given a priori, but we prepare them as a result of previous recursion steps. This results in a trade-off, converting circuits with exponential depth into those with exponential width. Nonetheless, QDP remains the strategically correct choice. Any protocol must conclude within the coherence time of the quantum processor, and QDP precisely offers an explicit tool that circumvents such hard limitations—providing a concrete strategy that radically reconfigures circuit structures to leverage quantum memoization.
@@
In this manuscript, we systematically introduce QDP for generic quantum recursions and establish its feasibility. We exemplify the utility of QDP on known algorithms, e.g. recursive Grover’s search [5] and quantum imaginary-time evolution [9]. Lastly, we propose a natively dynamic algorithm for transforming a state into its Schmidt basis without learning the state itself. With this, we establish that QDP is not only a strategic conceptual framework but also a highly applicable one.
Memory-calls in quantum recursions.—Quantum recursion unitaries can be regarded as a mapping from a quantum state $|{\psi}\rangle$ to a unitary operator $\hat{U}^{[\psi]}$. In general, $\hat{U}^{[\psi]}=e^{if(|{\psi}\rangle\!\langle{\psi}|)}$, where $f$ is a function from Hermitian operators to Hermitian operators. Then, $\hat{U}^{[\psi]}$ can be approximated by
$\hat{U}^{[\{\hat{\mathcal{N}}\},\psi]}=\hat{V}_{L}e^{i\hat{\mathcal{N}}_{L}(|{\psi}\rangle\!\langle{\psi}|)}\hat{V}_{L-1}\ldots\hat{V}_{1}e^{i\hat{\mathcal{N}}_{1}(|{\psi}\rangle\!\langle{\psi}|)}\hat{V}_{0}\ ,$ (2)
where $\{\hat{\mathcal{N}}\}$ is a collection of Hermitian-preserving (linear or polynomial) maps $\hat{\mathcal{N}}_{i}$, while $\hat{V}_{i}$ are static unitaries independent of the instruction $|{\psi}\rangle$. We say that a quantum recursion unitary written as Eq. (2) contains $L$ memory-calls of the form $e^{i\hat{\mathcal{N}}(|{\psi}\rangle\!\langle{\psi}|)}$. The most prominent example is found in Grover’s algorithm [25, 4], where alternating reflections
$\hat{U}^{[\psi]}_{L}=\displaystyle\prod_{i=1}^{L}e^{-i\alpha_{i}|{\psi}\rangle\!\langle{\psi}|}e^{-i\beta_{i}|{\tau}\rangle\!\langle{\tau}|}\ ,$ (3)
with respect to the initial state $|{\psi}\rangle$ and the target state of the search $|{\tau}\rangle$ are applied to $|{\psi}\rangle$. Eq. (3) contains $L$ memory-calls, each $e^{i\hat{\mathcal{N}}_{i}(\psi)}$ defined by $\hat{\mathcal{N}}_{i}(\psi)=-\alpha_{i}|{\psi}\rangle\!\langle{\psi}|$ and a set of predetermined angles $\{\alpha_{i},\beta_{i}\}_{i=1}^{L}$. Another example is the double-bracket iteration $\hat{U}^{[\psi]}=e^{s[\hat{D},|{\psi}\rangle\!\langle{\psi}|]}$ with a fixed operator $\hat{D}$, which facilitates diagonalization [6]. This corresponds to choosing $L=1$, $V_{0}=V_{1}=\mathmybb{1}$, and $\hat{\mathcal{N}}_{1}(|{\psi}\rangle\!\langle{\psi}|)=-is[\hat{D},|{\psi}\rangle\!\langle{\psi}|]$.
Unfolding quantum recursions.— We start by introducing the standard approach in implementing quantum recursions. Consider the Grover unitary Eq. (3), but applied recursively $|{\psi_{n+1}}\rangle=\hat{U}^{[\psi_{n}]}_{L}|{\psi_{n}}\rangle$ starting from a root state $|{\psi_{0}}\rangle$. By assumption, $e^{-i\alpha_{i}|{\psi_{0}}\rangle\!\langle{\psi_{0}}|}$ and $e^{-i\beta_{i}|{\tau}\rangle\!\langle{\tau}|}$ are readily available. However, $\hat{U}^{[\psi_{n}]}_{L}$ involves memory-calls $e^{-i\alpha_{i}|{\psi_{n}}\rangle\!\langle{\psi_{n}}|}$, which are not directly accessible. Using the relation $|{\psi_{n}}\rangle\!\langle{\psi_{n}}|=\hat{U}^{[\psi_{n-1}]}_{L}|{\psi_{n-1}}\rangle\!\langle{\psi_{n-1}}|\bigl{(}\hat{U}^{[\psi_{n-1}]}_{L}\bigr{)}^{\dagger}$, we observe that
$e^{-i\alpha_{i}|{\psi_{n}}\rangle\!\langle{\psi_{n}}|}=\hat{U}^{[\psi_{n-1}]}_{L}e^{-i\alpha_{i}|{\psi_{n-1}}\rangle\!\langle{\psi_{n-1}}|}\bigl{(}\hat{U}^{[\psi_{n}-1]}_{L}\bigr{)}^{\dagger}\ .$ (4)
We name this property covariance and use it to ‘unfold’ the memory-call $e^{-i\alpha_{i}|{\psi_{n}}\rangle\!\langle{\psi_{n}}|}$ into $12L+$ 1 memory-calls to $|{\psi_{n-1}}\rangle$. This unfolding continues until each memory-call to $|{\psi_{n}}\rangle$ in $\hat{U}^{[\psi_{n}]}_{L}$ is implemented with $(2L+1)^{n}$ calls to $|{\psi_{0}}\rangle$. The nested fixed-point Grover search [5] exactly follows this method.
To our knowledge, all existing quantum recursions are implemented through unfolding. However, there are two severe drawbacks. First, memory-calls must be covariant; otherwise they need to be approximated by covariant ones, via group commutators [26] as in Refs. [6, 7, 8, 9] or a general scheme described in Sec. II.1 of [27]. Moreover, unfolding yields exponentially deep circuits, with the final recursion step $\hat{U}^{[\psi_{N-1}]}$ dominating the depth, as it requires $L(2L+1)^{N-1}$ calls to $|{\psi_{0}}\rangle$.
Memory-usage queries.—To resolve the exponential depth of unfolded circuits, QDP invokes dynamically evolving memory states to directly instruct memory-calls. The goal is to apply the unitary operation $e^{i\hat{\mathcal{N}}(|{\psi}\rangle\!\langle{\psi}|)}$ to a working state $|{\phi}\rangle$:
$|{\phi}\rangle\mapsto e^{i\hat{\mathcal{N}}(|{\psi}\rangle\!\langle{\psi}|)}|{\phi}\rangle\ .$ (5)
A crucial observation is that Eq. (5) can be approximately implemented using multiple copies of the (unknown) state $|{\psi}\rangle$ [16, 18]. Consider the channel
^s$\hat{\mathcal{E}}^{[\hat{\mathcal{N}},\psi]}_{s}(|{\phi}\rangle\!\langle{\phi}|)=\operatorname{Tr}_{1}\left[e^{-i\hat{N}s}\left(|{\psi}\rangle\!\langle{\psi}|\otimes|{\phi}\rangle\!\langle{\phi}|\right)e^{i\hat{N}s}\right]\ ,$ (6)
@@
Exponential depth reduction with QDP.—We define QDP as the process of making memory-usage queries on copies of $|{\psi_{n}}\rangle$ to approximate a recursion step towards $|{\psi_{n+1}}\rangle$. Specifically, we initialize a quantum program with $\sigma_{0}=|{\psi_{0}}\rangle\!\langle{\psi_{0}}|$ and define the QDP iteration as
$\sigma_{n+1}=\hat{V}_{1}\left(\hat{\mathcal{E}}^{[\hat{\mathcal{N}},\sigma_{n}]}_{1/M}\right)^{M}\left(\hat{V}_{0}\sigma_{n}\hat{V}_{0}^{\dagger}\right)\hat{V}_{1}^{\dagger}\ .$ (7)
@@
Eq. (7) approximates a single memory-call recursion (Eq. (2) with $L=1$) by replacing the memory-call with $M$ memory-usage queries. In general, memory states $\sigma_{n\neq 0}$ are not pure because the channel ^$\hat{\mathcal{E}}^{[\hat{\mathcal{N}},\sigma_{n}]}_{1/M}$ is not exactly unitary. However, we show in Theorem 1 that $\sigma_{N}$ can be made arbitrarily close to the desired state $|{\psi_{N}}\rangle$. If recursions involve multiple memory-calls, QDP generalizes by replacing all memory-calls with $M$ total memory-usage queries to the memory state $\sigma_{n}$.
To prepare one copy of $\sigma_{1}$, we make $M$ memory-usage queries with the memory state $\sigma_{0}$. This requires $(M+1)$ root state copies ($M$ in the memory register and one in the working register). Likewise, preparing $\sigma_{2}$ requires $(M+1)^{2}$ copies of $\sigma_{0}$; thus, the iteration of Eq. (7) consumes $(M+1)^{n}$ copies of the root state for $\sigma_{n}$. Meanwhile, the circuit depth remains linear: multiple instruction states can be prepared in parallel, resulting in a maximum depth of $nM$ memory-usage queries. An intuitive explanation can be drawn by comparing QDP to unfolding, where a memory-call to $|{\psi_{n}}\rangle$ is executed by making $(2L+1)^{n}$ memory-calls to the root state. In contrast, QDP folds all root state calls into the memory state $\sigma_{n}$, allowing a memory-call to be implemented with a fixed-depth circuit.
The accuracy of QDP.—Seeking an exact solution to quantum recursions is unrealistic, as circuit compilation always entails inaccuracies. Hence, it is necessary to ensure that each memory-call approximation in QDP is sufficiently accurate to prepare the final state $\sigma_{N}$ within the same error threshold as unfolding. Setting $M=\mathcal{O}(1/\epsilon)$ in Eq. (7) ensures that the QDP prepared state $\sigma_{1}$ approximates the exact result $|{\psi_{1}}\rangle$ within an error $\|\sigma_{1}-|{\psi_{1}}\rangle\!\langle{\psi_{1}}|\|_{1}\leq\mathcal{O}(\epsilon)$. However, subsequent steps may amplify this error, as $\sigma_{1}\mapsto\sigma_{2}$ becomes instructed by $\sigma_{1}$, not $|{\psi_{1}}\rangle$. Indeed, the triangle inequality gives $\|\sigma_{2}-|{\psi_{2}}\rangle\!\langle{\psi_{2}}|\|_{1}\leq\mathcal{O}(M\epsilon)$; see Sec. II of [27] for the full analysis.
To address this, we identify sufficient conditions to prevent such destructive error propagation. It turns out that if the first few steps are sufficiently accurate, then later steps benefit from stabilization properties of typical quantum recursions. In particular, recursions with fixed-points (e.g. via the Polyak-Łojasiewicz inequality in gradient descent iterations [28] or asymptotic stability in time-discrete dynamical systems [29]) typically exhibit exponential convergence. For quantum recursions, this implies $\||{\psi_{N}}\rangle-|{\psi_{\infty}}\rangle\|_{1}\leq\alpha^{N}\||{\psi_{0}}\rangle-|{\psi_{\infty}}\rangle\|_{1}$, for some $\alpha<1$. However, this relies on three assumptions: i) a stable fixed-point $|{\psi_{\infty}}\rangle$ as $N\rightarrow\infty$, ii) uniqueness of the fixed-point, and that iii) the fixed-point is sufficiently strongly attracting. We term quantum recursions satisfying i,ii,iii) (or their generalization to mixed state recursions as in Sec. III of [27]) as exhibiting fast spectral convergence. Without iii), recursions become unstable, suggesting limited physical relevance, as practically achievable protocols must withstand small perturbations. In such cases, quantum computation would require infinite precision and resources for a successful operation.
It is worth noting that unfolding is also subject to similar stability constraints: if $\hat{U}^{[\psi_{0}]}$ is not compiled exactly, subsequent recursion unitaries $\hat{U}^{[\psi_{n}]}$, which rely on exponentially many applications of $\hat{U}^{[\psi_{0}]}$, become exponentially unstable. Thus, whether using QDP or unfolding, quantum recursions must exhibit fast spectral convergence; otherwise, the computational task is ill-conditioned. When this condition is met, QDP delivers an exponential depth reduction compared to unfolding:
Theorem 1 (Exponential circuit depth reduction).
Suppose a quantum recursion starting from $|{\psi_{0}}\rangle$ satisfies fast spectral convergence. Then QDP combined with a mixedness reduction protocol yields a final state $\sigma_{N}$ satisfying
$\|\sigma_{N}-|{\psi_{N}}\rangle\!\langle{\psi_{N}}|\|_{1}\leq\epsilon,$ (8)